### v4

# Sparse-PGD: A Unified Framework for Sparse Adversarial Perturbations Generation

Abstract
This work studies sparse adversarial perturbations, including both unstructured and structured ones. We propose a framework based on a white-box PGD-like attack method named Sparse-PGD to effectively and efficiently generate such perturbations. Furthermore, we combine Sparse-PGD with a black-box attack to comprehensively and more reliably evaluate the models’ robustness against unstructured and structured sparse adversarial perturbations. Moreover, the efficiency of Sparse-PGD enables us to conduct adversarial training to build robust models against various sparse perturbations. Extensive experiments demonstrate that our proposed attack algorithm exhibits strong performance in different scenarios. More importantly, compared with other robust models, our adversarially trained model demonstrates state-of-the-art robustness against various sparse attacks. Codes are available at https://github.com/CityU-MLO/sPGD.
I Introduction
Deep learning has been developing tremendously fast in the last decade. However, it is shown vulnerable to adversarial attacks: imperceivable adversarial perturbations [1, 2] could change the prediction of a model without altering the input’s semantic content, which poses great challenges in safety-critical systems. Among different kinds of adversarial perturbations, the ones bounded by $l_{\infty}$ or $l_{2}$ norms are mostly well-studied [3, 4, 5] and benchmarked [6], because these adversarial budgets, i.e., the sets of all allowable perturbations, are convex, which facilitates theoretical analyses and algorithm design. By contrast, we study sparse perturbations in this work, including both unstructured and structured ones which are bounded by $l_{0}$ norm and group $l_{0}$ norm, respectively. These perturbations are quite common in physical scenarios, including broken pixels in LED screens to fool object detection models and adversarial stickers on road signs to make an auto-driving system fail [7, 8, 9, 10, 11].
However, constructing such adversarial perturbations is challenging as the corresponding adversarial budget is non-convex. Therefore, gradient-based methods, such as projected gradient descent (PGD) [4], usually cannot efficiently obtain a strong adversarial perturbation. In this regard, existing methods to generate sparse perturbations [12, 13, 14, 15, 16, 17] either cannot control the $l_{0}$ norm or the group $l_{0}$ norm of perturbations or have prohibitively high computational complexity, which makes them inapplicable for adversarial training to obtain robust models against sparse perturbations. The perturbations bounded by $l_{1}$ norm are the closest scenario to $l_{0}$ bounded perturbations among convex adversarial budgets defined by an $l_{p}$ norm. Nevertheless, adversarial training in this case [18, 19] still suffers from issues such as slow convergence and instability. Jiang et al. [20] demonstrates that these issues arise from non-sparse perturbations bounded by $l_{1}$ norm. In other words, $l_{1}$ adversarial budget cannot guarantee the sparsity of the perturbations. Furthermore, existing works investigating structured sparse perturbations [21, 22, 23, 16, 17] only support generating a single adversarial patch, lacking flexibility and generality. Thus, it is necessary but challenging to develop a unified framework for both unstructured and structured sparse adversarial perturbations.
In this work, we propose a white-box attack named Sparse-PGD (sPGD) to effectively and efficiently generate unstructured and structured sparse perturbations. For unstructured sparse perturbations, we decompose the perturbation $\bm{\delta}$ as the product of a magnitude tensor $\bm{p}$} and a binary sparse mask $\bm{m}$}: $\bm{\delta}={\bm{p}}\odot{\bm{m}}$ , where $\bm{p}$} and $\bm{m}$} determine the magnitudes and the locations of perturbed features, respectively. Although $\bm{p}$} can be updated by PGD-like methods, it is challenging to directly optimize the binary mask $\bm{m}$} in the discrete space. We thereby introduce an alternative continuous variable $\widetilde{{\bm{m}}}$ to approximate $\bm{m}$} and update $\widetilde{{\bm{m}}}$ by gradient-based methods, $\widetilde{{\bm{m}}}$ is then transformed to $\bm{m}$} by projection to the discrete space. To further boost the performance, we propose the unprojected gradient of $\bm{p}$} and random reinitialization mechanism. For structured sparse perturbations, we study group $l_{0}$ norm and its approximated version based on the group norm proposed in [24]. The structured sparse perturbation bounded by approximated group $l_{0}$ norm can be thus decomposed as ${\bm{\delta}}={\bm{p}}\odot{\bm{m}}={\bm{p}}\odot\min\left(\mathrm{TConv}({\bm{v}},{\bm{k}}),1\right)$, where the binary group mask $\bm{v}$} determines the positions of the groups to be perturbed and the customized binary kernel $\bm{k}$} determines the pattern of groups. That is to say, we decompose the structured sparse perturbations into their positions and the patterns. Furthermore, we leverage transposed convolution ($\mathrm{TConv}$) and clipping operations to map the positions of the groups $\bm{v}$} to the positions of perturbed features $\bm{m}$}. Similar to the unstructured cases, we introduce the continuous $\widetilde{{\bm{v}}}$ and update it by gradient-based methods. Ultimately, we manage to transform the problem of optimizing the structured sparse mask $\bm{m}$} into the problem of optimizing the group mask $\bm{v}$} bounded by $l_{0}$ norm, which can be resolved in the framework of sPGD. The whole pipeline is illustrated in Figure 2. On top of sPGD, we propose Sparse-AutoAttack (sAA), which is the ensemble of the white-box sPGD and another black-box sparse attack, for a more comprehensive and reliable evaluation against both unstructured and structured sparse perturbations. Through extensive experiments, we show that our method exhibits better performance than other attacks.
More importantly, we explore adversarial training to obtain robust models against sparse attacks. In this context, the attack method will be called in each mini-batch update, so it should be both effective and efficient. Compared with existing methods, our proposed sPGD performs much better when using a small number of iterations, making it feasible for adversarial training and its variants [25]. Empirically, models adversarially trained by sPGD demonstrate the strongest robustness against various sparse attacks.
We summarize the contributions of this paper as follows:
  1. 1.
    We propose an effective and efficient white-box attack algorithm named Sparse-PGD (sPGD) which can be utilized to generate both unstructured and structured sparse adversarial perturbations.
  2. 2.
    sPGD achieves the best performance among white-box sparse attacks. We then combine it with a black-box sparse attack to construct Sparse-AutoAttack (sAA) for more comprehensive robustness evaluation against sparse adversarial perturbations.
  3. 3.
    sPGD achieves much better performance in the regime of limited iterations, it is then adopted for adversarial training. Extensive experiments demonstrate that models adversarially trained by sPGD have significantly stronger robustness against various sparse attacks.
Preliminaries: We use image classification as an example, although the proposed methods are applicable to any classification model. Under $l_{p}$ bounded perturbations, the robust learning aims to solve the following min-max optimization problem.
[…]
where $\bm{\theta}$ denotes the parameters of the model and $\mathcal{L}$} is the loss objective function. ${\bm{x}}_{i}\in\mathbb{R}^{h\times w\times c}$ is the input image where $h$, $w$ and $c$ represent the height, width, and number of channels, respectively. $\bm{\delta}_{i}$ has the same shape as ${\bm{x}}_{i}$ and represents the perturbation. The perturbations are constrained by its $l_{p}$ norm and the bounding box. In this regard, we use the term adversarial budget to represent the set of all allowable perturbations. Adversarial attacks focus on the inner maximization problem of (1) and aim to find the optimal adversarial perturbation, while adversarial training focuses on the outer minimization problem of (1) and aims to find a robust model parameterized by $\bm{\theta}$. Due to the high dimensionality and non-convexity of the loss function when training a deep neural network, [26] has proven that solving the problem (1) is at least NP-complete.
We consider the pixel sparsity for image inputs in this work, which is more meaningful than feature sparsity and consistent with existing works [13, 17]. That is, a pixel is considered perturbed if any of its channel is perturbed, and sparse perturbation means few pixels are perturbed.
II Related Works
Non-sparse Attacks: The pioneering work [1] finds the adversarial perturbations to fool image classifiers and proposes a method to minimize the $l_{2}$ norm of such perturbations. To more efficiently generate adversarial perturbations, the fast gradient sign method (FGSM) [3] generates $l_{\infty}$ perturbation in one step, but its performance is significantly surpassed by the multi-step variants [27]. Projected Gradient Descent (PGD) [4] further boosts the attack performance by using iterative updating and random initialization. Specifically, each iteration of PGD updates the adversarial perturbation $\bm{\delta}$ by:
$\bm{\delta}\longleftarrow\Pi_{{\mathcal{S}}}(\bm{\delta}+\alpha\cdot s(\nabla_{\bm{\delta}}\mathcal{L}({\bm{\theta}},{\bm{x}}+\bm{\delta})))$ (2)
where $\mathcal{S}$} is the adversarial budget, $\alpha$ is the step size, $s:\mathbb{R}^{h\times w\times c}\rightarrow\mathbb{R}^{h\times w\times c}$ selects the steepest ascent direction based on the gradient of the loss $\mathcal{L}$ with respect to the perturbation. Inspired by the first-order Taylor expansion, Madry et al. [4] derives the steepest ascent direction for $l_{2}$ bounded and $l_{\infty}$ bounded perturbations to efficiently find strong adversarial examples; SLIDE [18] and $l_{1}$-APGD [19] use $k$-coordinate ascent to construct $l_{1}$ bounded perturbations, which is shown to suffer from the slow convergence [20].
Besides the attacks that have access to the gradient of the input (i.e., white-box attacks), there are black-box attacks that do not have access to model parameters, including the ones based on gradient estimation through finite differences [28, 29, 30, 31, 32] and the ones based on evolutionary strategies or random search [33, 34]. To improve the query efficiency of these attacks, [35, 36, 37, 38] generate adversarial perturbation at the corners of the adversarial budget.
To more reliably evaluate the robustness, Croce and Hein [39] proposes AutoAttack (AA) which consists of an ensemble of several attack methods, including both black-box and white-box attacks. Croce and Hein [19] extends AA to the case of $l_{1}$ bounded perturbations and proposes AutoAttack-$l_{1}$ (AA-$l_{1}$). Although the $l_{1}$ bounded perturbations are usually sparse, Jiang et al. [20] demonstrates that AA-$l_{1}$ is able to find non-sparse perturbations that cannot be found by SLIDE to fool the models. That is to say, $l_{1}$ bounded adversarial perturbations are not guaranteed to be sparse. We should study perturbations bounded by $l_{0}$ norm.
Sparse Attacks: For perturbations bounded by $l_{0}$ norm, directly adopting vanilla PGD as in Eq. (2) leads to suboptimal performance due to the non-convexity nature of the adversarial budget: $\mathrm{PGD_{0}}$ [13], which updates the perturbation by gradient ascent and project it back to the adversarial budget, turns out very likely to trap in the local maxima. Different from $\mathrm{PGD_{0}}$, CW L0 [40] projects the perturbation onto the feasible set based on the absolute product of gradient and perturbation and adopts a mechanism similar to CW L2 [40] to update the perturbation. SparseFool [12] and GreedyFool [15] also generate sparse perturbations, but they do not strictly restrict the $l_{0}$ norm of perturbations. If we project their generated perturbations to the desired $l_{0}$ ball, their performance will drastically drop. Sparse Adversarial and Interpretable Attack Framework (SAIF) [41] is similar to our method in that SAIF also decomposes the $l_{0}$ perturbation into a magnitude tensor and sparsity mask, but it uses the Frank-Wolfe algorithm [42] to separately update them. SAIF turns out to get trapped in local minima and shows poor performance on adversarially trained models. Besides white-box attacks, there are black-box attacks to generate sparse adversarial perturbations, including CornerSearch [13] and Sparse-RS [17]. However, these black-box attacks usually require thousands of queries to find an adversarial example, making it difficult to scale up to large datasets. In addition to the unstructured adversarial perturbations mentioned above, there are several works discussing structured sparse perturbations, including universal adversarial patch for all data [21, 22] and some image-specific patches [23, 16, 17]. However, these works only support generating a single adversarial patch, which lacks of flexibility and generality.
Adversarial Training: Despite the difficulty in obtaining robust deep neural network, adversarial training [4, 43, 44, 45, 46, 47, 48, 49] stands out as a reliable and popular approach to do so [50, 39]. It generates adversarial examples first and then uses them to optimize model parameters. Despite effective, adversarial training is time-consuming due to multi-step attacks. Shafahi et al. [51], Zhang et al. [52], Wong et al. [53], Sriramanan et al. [54] use weaker but faster one-step attacks to reduce the overhead, but they may suffer from catastrophic overfitting [55]: the model overfits to these weak attacks during training instead of achieving true robustness to various attacks. Kim et al. [56], Andriushchenko and Flammarion [57], Golgooni et al. [58], de Jorge et al. [59] try to overcome catastrophic overfitting while maintaining efficiency.
Compared with $l_{\infty}$ and $l_{2}$ bounded perturbations, adversarial training against $l_{1}$ bounded perturbations is shown to be even more time-consuming to achieve the optimal performance [19]. In the case of $l_{0}$ bounded perturbations, PGD0 [13] is adopted for adversarial training. However, models trained by PGD0 exhibit poor robustness against strong sparse attacks. In this work, we propose an effective and efficient sparse attack that enables us to train a model that is more robust against various sparse attacks than existing methods.
III Unstructured Sparse Adversarial Attack
In this section, we introduce Sparse-PGD (sPGD) for generating unstructured sparse perturbations. Its extension that generates structured sparse perturbations is introduced in Sec. IV. Similar to AutoAttack [39, 19], we further combine sPGD with a black-box attack to construct sparse-AutoAttack (sAA) for more comprehensive and reliable robustness evaluation.
III-A Sparse-PGD (sPGD)
Inspired by SAIF [41], we decompose the sparse perturbation $\bm{\delta}$ into a magnitude tensor ${\bm{p}}\in\mathbb{R}^{h\times w\times c}$ and a sparsity mask ${\bm{m}}\in\{0,1\}^{h\times w\times 1}$, i.e., $\bm{\delta}={\bm{p}}\odot{\bm{m}}$. Therefore, the attacker aims to maximize the following loss objective function:
$\max_{\|\bm{\delta}\|_{0}\leq\epsilon,0\leq{\bm{x}}+\bm{\delta}\leq 1}{\mathcal{L}}(\bm{\theta},{\bm{x}}+\bm{\delta})=\max_{{\bm{p}}\in{\mathcal{S}}_{\bm{p}},{\bm{m}}\in{\mathcal{S}}_{\bm{m}}}\mathcal{L}(\bm{\theta},{\bm{x}}+{\bm{p}}\odot{\bm{m}}).$ (3)
The feasible sets for $\bm{p}$} and $\bm{m}$} are ${\mathcal{S}}_{\bm{p}}=\{{\bm{p}}\in\mathbb{R}^{h\times w\times c}|0\leq{\bm{x}}+{\bm{p}}\leq 1\}$ and ${\mathcal{S}}_{\bm{m}}=\{{\bm{m}}\in\{0,1\}^{h\times w\times 1}|\|{\bm{m}}\|_{0}\leq\epsilon\}$, respectively. Similar to PGD, sPGD iteratively updates $\bm{p}$} and $\bm{m}$} until finding a successful adversarial example or reaching the maximum iteration number.
Update Magnitude Tensor $\bm{p}$}: The magnitude tensor $\bm{p}$} is only constrained by the input domain. In the case of images, the input is bounded between 0 and $1$. Note that the constraints on $\bm{p}$} are elementwise and similar to those of $l_{\infty}$ bounded perturbations. Therefore, instead of greedy or random search [13, 17], we utilize PGD in the $l_{\infty}$ case, i.e., use the sign of the gradients, to optimize $\bm{p}$} as demonstrated by Eq. (4) below, with $\alpha$ being the step size.
${\bm{p}}\longleftarrow\Pi_{{\mathcal{S}}_{{\bm{p}}}}\left({\bm{p}}+\alpha\cdot\mathtt{sign}(\nabla_{{\bm{p}}}{\mathcal{L}}(\bm{\theta},{\bm{x}}+{\bm{p}}\odot{\bm{m}}))\right),$ (4)
Update Sparsity Mask $\bm{m}$}: The sparsity mask $\bm{m}$} is binary and constrained by its $l_{0}$ norm. Directly optimizing the discrete variable $\bm{m}$} is challenging, so we update its continuous alternative $\widetilde{{\bm{m}}}\in\mathbb{R}^{h\times w\times 1}$ and project $\widetilde{{\bm{m}}}$ to the feasible set ${\mathcal{S}}_{\bm{m}}$ to obtain $\bm{m}$} before multiplying it with the magnitude tensor $\bm{p}$} to obtain the sparse perturbation $\bm{\delta}$. Specifically, $\widetilde{{\bm{m}}}$ is updated by gradient ascent. Projecting $\widetilde{{\bm{m}}}$ to the feasible set ${\mathcal{S}}_{{\bm{m}}}$ is to set the $\epsilon$-largest elements in $\widetilde{{\bm{m}}}$ to $1$ and the rest to 0. In addition, we adopt the sigmoid function to normalize the elements of $\widetilde{{\bm{m}}}$ before projection.
Mathematically, the update rules for $\widetilde{{\bm{m}}}$ and $\bm{m}$} are demonstrated as follows:
$\widetilde{{\bm{m}}}$ $\longleftarrow\widetilde{{\bm{m}}}+\beta\cdot\nabla_{\widetilde{{\bm{m}}}}{\mathcal{L}}/\|\nabla_{\widetilde{{\bm{m}}}}{\mathcal{L}}\|_{2},$ (5)
${\bm{m}}$ $)\displaystyle\longleftarrow\Pi_{{\mathcal{S}}_{{\bm{m}}}}(\sigma(\widetilde{{\bm{m}}}$)) (6)
where $\beta$ is the step size for updating the sparsity mask’s continuous alternative $\widetilde{{\bm{m}}}$, $\sigma(\cdot)$ denotes the sigmoid function. Furthermore, to prevent the magnitude of $\widetilde{{\bm{m}}}$ from becoming explosively large, we do not update $\widetilde{{\bm{m}}}$ when $\nabla_{\widetilde{{\bm{m}}}}{\mathcal{L}}||_{2}<\gamma$, which indicates that $\widetilde{{\bm{m}}}$ is located in the saturation zone of sigmoid function. The gradient $\nabla_{\widetilde{{\bm{m}}}}{\mathcal{L}}$ is calculated at the point $)\bm{\delta}={\bm{p}}\odot\Pi_{{\mathcal{S}}_{{\bm{m}}}}(\sigma(\widetilde{{\bm{m}}}$)), where the loss function is not always differentiable. We demonstrate how to estimate the update direction in the next part.
[…]
In practice, the performance of using $g_{{\bm{p}}}$ and $\widetilde{g}_{{\bm{p}}}$ to optimize $\bm{p}$} is complementary. The sparse gradient $g_{{\bm{p}}}$ is consistent with the forward propagation and is thus better at exploitation. By contrast, the unprojected gradient $\widetilde{g}_{{\bm{p}}}$ updates the $\bm{p}$} by a dense tensor and is thus better at exploration. In view of this, we set up an ensemble of attacks with both gradients to balance exploration and exploitation.
[…]
Algorithm 1 Sparse-PGD for $l_{0}$ Bounded Perturbations
1:Input: Clean image: ${\bm{x}}\in[0,1]^{h\times w\times c}$; Model parameters: $\bm{\theta}$}; Max iteration number: $T$; Tolerance: $t$; $l_{0}$ budget: $\epsilon$; Step size: $\alpha$, $\beta$; Small constant: $\gamma=2\times 10^{-8}$
2: Random initialize $\bm{p}$} and $\widetilde{{\bm{m}}}$
[…]
4:for i=0,1,,T1i=0,1,...,T-1 do
5:  Calculate the loss $\mathcal{L}({\bm{\theta}},{\bm{x}}+{\bm{p}}\odot{\bm{m}})$
6:  if unprojected then
[…]
8:  else
9:   $g_{{\bm{p}}}=\nabla_{\bm{\delta}}{\mathcal{L}}\odot{\bm{m}}$
10:  end if
[…]
12:  ${\bm{p}}=\Pi_{{\mathcal{S}}_{{\bm{p}}}}({\bm{p}}+\alpha\cdot\mathtt{sign}(g_{{\bm{p}}}))$
13:  ${\bm{d}}=g_{\widetilde{{\bm{m}}}}/\|g_{\widetilde{{\bm{m}}}}\|_{2}$ if $g_{\widetilde{{\bm{m}}}}||_{2}\geq\gamma$ else 0
14:  ${\bm{m}}_{old},\ \widetilde{{\bm{m}}}={\bm{m}},\ \widetilde{{\bm{m}}}+\beta\cdot{\bm{d}}$
15:  $){\bm{m}}=\Pi_{{\mathcal{S}}_{{\bm{m}}}}(\sigma(\widetilde{{\bm{m}}}$))
16:  if attack succeeds then
17:   break
18:  end if
19:  if ${\bm{m}}-{\bm{m}}_{old}||_{0}\leq 0$ for $t$ consecutive iters then
20:   Random initialize $\widetilde{{\bm{m}}}$
21:  end if
22:end for
23:Output: Perturbation: $\bm{\delta}={\bm{p}}\odot{\bm{m}}$
Random Reinitialization: Due to the projection to the set ${\mathcal{S}}_{{\bm{m}}}$ in Eq. (6), the sparsity mask $\bm{m}$} changes only when the relative magnitude ordering of the continuous alternative $\widetilde{{\bm{m}}}$ changes. In other words, slight changes in $\widetilde{{\bm{m}}}$ usually mean no change in $\bm{m}$}. As a result, $\bm{m}$} usually gets trapped in a local maximum. To solve this problem, we propose a random reinitialization mechanism. Specifically, when the attack fails, i.e., the model still gives the correct prediction, and the current sparsity mask $\bm{m}$} remains unchanged for three consecutive iterations, the continuous alternative $\widetilde{{\bm{m}}}$ will be randomly reinitialized for better exploration.
To summarize, we provide the pseudo-code of sparse PGD (sPGD) in Algorithm 1. SAIF [41] also decomposes the perturbation $\bm{\delta}$ into a magnitude tensor $\bm{p}$} and a mask $\bm{m}$}, but uses a different update rule: it uses Frank-Wolfe to update both $\bm{p}$} and $\bm{m}$}. By contrast, we introduce the continuous alternative $\widetilde{{\bm{m}}}$ of $\bm{m}$} and use gradient ascent to update $\bm{p}$} and $\widetilde{{\bm{m}}}$. Moreover, we include unprojected gradient and random reinitialization techniques in Algorithm 1 to further enhance the performance.
III-B Sparse-AutoAttack (sAA)
AutoAttack (AA) [39] is an ensemble of four diverse attacks for a standardized parameters-free and reliable evaluation of robustness against $l_{\infty}$ and $l_{2}$ attacks. Croce and Hein [19] extends AutoAttack to $l_{1}$ bounded perturbations. In this work, we propose sparse-AutoAttack (sAA), which is also a parameter-free ensemble of both black-box and white-box attacks for comprehensive robustness evaluation against $l_{0}$ bounded perturbations. It can be used in a plug-and-play manner. However, different from the $l_{\infty}$, $l_{2}$ and $l_{1}$ cases, the adaptive step size, momentum and difference of logits ratio (DLR) loss function do not improve the performance in the $l_{0}$ case, so they are not adopted in sAA. In addition, compared with targeted attacks, sPGD turns out stronger when using a larger query budget in the untargeted settings given the same total number of back-propagations. As a result, we only include the untargeted sPGD with cross-entropy loss and constant step sizes in sAA. Specifically, we run sPGD twice for two different backward functions: one denoted as sPGDproj uses the sparse gradient $g_{{\bm{p}}}$, and the other denoted as sPGDunproj uses the unprojected gradient $\widetilde{g}_{{\bm{p}}}$ as described in Section III-A. As for the black-box attack, we adopt the strong black-box attack Sparse-RS [17], which can generate $l_{0}$ bounded perturbations. We run each version of sPGD and Sparse-RS for $10000$ iterations, respectively. We use cascade evaluation to improve the efficiency. Concretely, suppose we find a successful adversarial perturbation by one attack for one instance. Then, we will consider the model non-robust in this instance and the same instance will not be further evaluated by other attacks. Based on the efficiency and the attack success rate, the attacks in sAA are sorted in the order of sPGDunproj, sPGDproj and Sparse-RS.
IV Structured Sparse Adversarial Attack
In this section, we extend Sparse-PGD (sPGD) and Sparse-AutoAttack (sAA) to generate structured sparse perturbations.
IV-A Formulation of Structured Sparsity
Given an input ${\bm{x}}\in\mathbb{R}^{d}$, we can partition its $d$ features into several groups that may overlap and define structured sparsity based on them. These groups can represent pixels of a row, a column, a patch, or a particular pattern. Without the loss of generality, we use $\{1,2,...,d\}$ as the indices of $d$ features of the input and $N$ set of indices ${\mathcal{G}}=\{G_{j}\}_{j=1}^{N}$ to represent the groups. For any perturbation, we define its group $l_{0}$ norm based on groups $\mathcal{G}$} as the minimal number of groups needed to cover its non-zero components. Mathematically, Bach [24] proposed a convex envelope of group $l_{0}$ norm as follows:
$\Omega({\bm{x}})=\sum_{k=1}^{d}|x_{\pi_{k}}|\left[F(\{\pi_{1},...,\pi_{k}\})-F(\{\pi_{1},...,\pi_{k-1}\})\right].$ (7)
where $F$ is a submodular function defined on the subsets of $V=\{1,2,...,d\}$ and $\{\pi_{i}\}_{i=1}^{d}$ is a permutation of $\{1,2,...,d\}$. More specifically, $F(A)$ indicates the minimal number of groups from $\mathcal{G}$} to cover the set $A$ and $\{\pi_{i}\}_{i=1}^{d}$ indicates the components of $|{\bm{x}}|$ in the decreasing order, i.e., $|x_{\pi_{1}}|\geq...\geq|x_{\pi_{d}}|\geq 0$. Compared with the convex envelope $\Omega$, the group $l_{0}$ norm $\Omega_{0}$ is defined as follows:
$\Omega_{0}({\bm{x}})$ $=\sum_{k=1}^{d}\mathbbm{1}(|x_{\pi_{k}}|)\left[F(\{\pi_{1},...,\pi_{k}\})-F(\{\pi_{1},...,\pi_{k-1}\})\right]$ (8)
$=F(\{\pi_{1},.,\pi_{d^{\prime}}\})\ \ \mathrm{where}\ x_{\pi_{d^{\prime}}}\neq 0\ \mathrm{and}\ x_{\pi_{d^{\prime}+1}}=0.$
Similar to the difference between $l_{1}$ norm and $l_{0}$ norm, we replace $|x_{\pi_{k}}|$ with $\mathbbm{1}(|x_{\pi_{k}}|)$ in Eq. (8) to indicate the number of groups with non-zero entities, where $\mathbbm{1}$ is an indicator function. The second equality in Eq. (8) is based on the decreasing order of $\{|x_{\pi_{i}}|\}_{i=1}^{d}$ and is a more straightforward definition of the group $l_{0}$ norm. Like the $l_{0}$ norm, the adversarial budgets based on the group $l_{0}$ norm are not convex, either.
In general, it is difficult to decide the minimum number of groups to cover non-zero elements, i.e., to calculate the function $F$. Therefore, it is challenging to directly apply sPGD to the group $l_{0}$ norm constraints. To address this issue, we can approximate $\Omega_{0}$ by a tight surrogate that facilitates the optimization by sPGD. In this regard, we introduce a binary group mask ${\bm{v}}=[v_{1},...,v_{N}]\in\{0,1\}^{N}$ to indicate whether a specific group is chosen to be perturbed. Mathematically, $\forall i$, if ${\bm{x}}_{i}\neq 0$, then $\exists j\in\{1,2,...,N\}$, $v_{j}=1$ and $i\in G_{j}$. In this context, we propose the following approximation of the group $l_{0}$ norm as the surrogate:
$\Omega_{0}^{\prime}({\bm{x}},{\bm{v}})=\sum_{i=1}^{N}\mathbbm{1}(\|{\bm{x}}_{G_{i}}\|_{p})\cdot v_{i},$ (9)
where ${\bm{x}}_{G_{i}}$ is the subvector of $\bm{x}$} on the indices in $G_{i}$, and $p\in[0,+\infty]$. Compared with $\Omega_{0}$ in Eq. (8), $\Omega^{\prime}_{0}$ in Eq. (9) does not require the number of groups in $\mathcal{G}$} to cover non-zero elements in $\bm{x}$} to be minimum. Instead, $\Omega^{\prime}_{0}$ calculates the number of perturbed groups in $\mathcal{G}$} selected by the binary vector $\bm{v}$}. In addition, when $\mathbbm{1}(\|{\bm{x}}_{G_{i}}\|_{p})=1$ for all $i\in\{j~|~v_{j}=1\}$, the approximated group $l_{0}$ norm $\Omega_{0}^{\prime}({\bm{x}},{\bm{v}})$ is equivalent to the $l_{0}$ norm of the group mask $\bm{v}$}, i.e., $\Omega_{0}^{\prime}({\bm{x}},{\bm{v}})=\|{\bm{v}}\|_{0}$. In summary, we have the following theorem:
Theorem IV.1.
Given an input ${\bm{x}}\in\mathbb{R}^{d}$, a set of groups $\mathcal{G}=\{G_{j}\}_{j=1}^{N}$ and any vector $\bm{v}$} satisfying the constraint in the definition of $\Omega^{\prime}_{0}$, then we have $\Omega_{0}({\bm{x}})\leq\Omega_{0}^{\prime}({\bm{x}},{\bm{v}})\leq\|{\bm{v}}\|_{0}$.
Proof.
The first inequality is based on the minimum optimality condition of the function $F$ in Eq. (8). Since $\Omega_{0}({\bm{x}})$ indicates the minimal number of groups needed to cover the indices of non-zero elements in $\bm{x}$}, we have at least the same number of terms in Eq. (9) where ${\bm{x}}_{G_{i}}\neq 0$ and $v_{i}=1$. Therefore, we have $\Omega_{0}({\bm{x}})\leq\Omega^{\prime}_{0}({\bm{x}})$. For the second inequality, $\bm{v}$} is a binary vector, i.e., $\forall i$, ${\bm{v}}_{i}\in\{0,1\}$, so $\Omega_{0}^{\prime}({\bm{x}},{\bm{v}})\leq\|{\bm{v}}\|_{0}$ is clear.
Theorem IV.1 indicates that the number of perturbed groups determined by $\bm{v}$} can be larger than the minimal number of groups to cover the perturbed features $\bm{x}$}, i.e., $\Omega_{0}^{\prime}({\bm{x}},{\bm{v}})$ can be larger than $\Omega_{0}({\bm{x}})$. The gap between $\Omega_{0}^{\prime}({\bm{x}},{\bm{v}})$ and $\Omega_{0}({\bm{x}})$ stem from the potential overlap among groups in $\mathcal{G}$}. For instance, if we generate $1\times 3$ perturbations, for a $\bm{x}$} with the shape of $1\times 5$ and all of its entities are perturbed, the corresponding $\Omega_{0}({\bm{x}})=2$, while the possible $\|{\bm{v}}\|_{0}$ can be $2$ or $3$, i.e., ${\bm{v}}=[1,0,1]$ or $[1,1,1]$. Thus, we have $\Omega_{0}({\bm{x}})\leq\Omega_{0}^{\prime}({\bm{x}},{\bm{v}})\leq\|{\bm{v}}\|_{0}$. However, we demonstrate that $\Omega_{0}^{\prime}({\bm{\delta}},{\bm{v}})$ is quite close to $\Omega_{0}({\bm{\delta}})$ in practice. To corroborate this, we calculate the ratio between the group $l_{0}$ norm and approximated group $l_{0}$ norm, i.e., $\Omega_{0}({\bm{\delta}})/\Omega_{0}^{\prime}({\bm{\delta}},{\bm{v}})$ of perturbations under different sparsity levels $\epsilon$ defined by $\Omega_{0}^{\prime}({\bm{\delta}},{\bm{v}})\leq\epsilon$. As illustrated in Figure 1, when applying patch perturbations, we can observe that the ratio $\Omega_{0}({\bm{\delta}})/\Omega_{0}^{\prime}({\bm{\delta}},{\bm{v}})$ increases as the ratio between the maximum number of perturbed pixels and the total number of pixels decreases, because the probability of overlapping accordingly declines. Specifically, when $\epsilon$ is smaller than $5$, the ratios of perturbations on both CIFAR-10 and ImageNet-100 are larger than $0.95$. Notably, on ImageNet-100, the $l_{0}$ norm of $5$ non-overlapping $10\times 10$ patches reaches $500$, which is very large for a sparse perturbation used in practice, so we can conclude based on the simulation results that $\Omega_{0}^{\prime}$ is a good approximation of $\Omega_{0}$ across datasets with different resolutions.
Figure 1: The ratio between the group $l_{0}$ norm and the approximated group $l_{0}$ norm. The x-axis is the approximated group $l_{0}$ norm ranging from 1 to 25, and the y-axis is the ratio. We plot the ratios of $3\times 3$ patch perturbations on CIFAR-10 ($32\times 32$) and $10\times 10$ patch perturbations on ImageNet-100 ($224\times 224$), respectively. The results are calculated on 100 samples. The solid line and shadow denote the mean value and standard deviation, respectively.
IV-B sPGD for Structured Sparse Perturbations
Figure 2: Pipeline of sPGD for structured sparse perturbations. The continuous group mask $\widetilde{{\bm{v}}}$ is first projected to get the binary group mask $\bm{v}$} to ensure $\|{\bm{v}}\|_{0}\leq\epsilon$, which is similar to Eq. (6). Given the kernel ${\bm{k}}\in\{0,1\}^{r\times r}$ with a customized pattern, we can transform $\bm{v}$} to the pixel mask $\bm{m}$} using transposed convolution and clipping (see Eq. (10)). Finally, we element-wisely multiply $\bm{m}$} with the dense perturbation $\bm{p}$} to obtain the structured sparse perturbation $\bm{\delta}$}. Note that the continuous group mask $\widetilde{{\bm{v}}}$ and the dense perturbation $\bm{p}$} are learnable.
Based on the approximated group $l_{0}$ norm $\Omega_{0}^{\prime}$ proposed in Eq. (9), we extend Sparse-PGD (sPGD) to generate structured sparse perturbations. Similarly, we decompose the perturbation $\bm{\delta}$} into a magnitude tensor $\bm{p}$} and a binary pixel mask $\bm{m}$}. Like unstructured cases, the magnitude tensor $\bm{p}$} is updated using $l_{\infty}$ PGD, as in Eq. (4).
Based on the definition of structured sparsity and $\Omega^{\prime}_{0}$ defined in Eq. (9), if a specific group $G_{i}$ is chosen to be perturbed, then we let $v_{i}=1$, otherwise $v_{i}$ will be set 0. In this regard, finding the optimal $\bm{m}$} s.t. $\Omega^{\prime}_{0}({\bm{m}},{\bm{v}})\leq\epsilon$ is equivalent to finding the optimal group mask $\bm{v}$} s.t. $\|{\bm{v}}\|_{0}\leq\epsilon$, which can be resolved by sPGD. Similar to the unstructured cases, we optimize the continuous alternative $\widetilde{{\bm{v}}}$ rather than optimizing a binary $\bm{v}$} directly. In the following, we will elaborate on the mapping from $\bm{v}$} to $\bm{m}$} using different examples.
Column / Row: If we aim to perturb at most $\epsilon$ columns in an image ${\bm{x}}\in[0,1]^{h\times w\times c}$. Since we can partition $\bm{x}$} into $w$ groups (i.e., columns), $\bm{v}$} has the size of $1\times w$. Thus, $\bm{m}$} can obtained through expanding $\bm{v}$} from $1\times w$ to $h\times w$. As for row, the only difference is that the size of $\bm{v}$} is $h\times 1$. Note that, there is no overlap between any two columns or rows, so $\Omega^{\prime}_{0}({\bm{m}},{\bm{v}})\equiv\Omega_{0}({\bm{m}})$ in this case.
Patch / Any Pattern: Without loss of generality, we assume that the perturbations are contained to be localized in $r\times r$ patches. To avoid wasting budget due to the potential overlap at the border or corner of images, we let ${\bm{v}}\in\{0,1\}^{(h-r+1)\times(w-r+1)}$, and we can derive the mapping operation from $\bm{v}$} to $\bm{m}$} as follows:
${\bm{m}}=\min\left(\mathrm{TConv}({\bm{v}},{\bm{k}},s),1\right),$ (10)
where $\mathrm{TConv}$ is the transposed convolution operation, ${\bm{k}}=\mathbf{1}^{r\times r}$ is the kernel with all-one entities, $s=1$ denotes the stride. The clipping operation is to ensure the output value assigned to $\bm{m}$} is binary. Moreover, we can customize the kernel ${\bm{k}}\in\{0,1\}^{r\times r}$ in Eq. (10) to make the perturbations localized in groups with any desired patterns, such as hearts, stars, letters and so on. Notably, when ${\bm{k}}=[1]$, i.e., a $1\times 1$ matrix with a sole value $1$, the resulting perturbations are degraded to unstructured sparse ones; when ${\bm{k}}=\mathbf{1}^{h\times 1}$ or $\mathbf{1}^{1\times w}$, specific columns or rows are perturbed. The pipeline of sPGD for structured sparse perturbations in the general case is illustrated in Figure 2.
In addition to the forward mapping operation defined in Eq. (10), we provide its backward function as follows:
${\bm{g}}_{{\bm{v}}}=\mathrm{Conv}({\bm{g}}_{{\bm{m}}},{\bm{k}},s),$ (11)
where $\mathrm{Conv}$ is the convolution operation, and ${\bm{g}}_{{\bm{m}}}$ is the gradient of $\bm{m}$}. We neglect the clipping operation when calculating the gradient of $\bm{v}$} since it is not always differentiable.
The extension of sPGD for structured sparse perturbations offers a unified framework to generate sparse adversarial perturbations. We provide its pseudo-code in Algorithm 2.
Algorithm 2 Sparse-PGD for Structured Sparse Perturbations
1:Input: Clean image: ${\bm{x}}\in[0,1]^{h\times w\times c}$; Model parameters: $\bm{\theta}$}; Customized pattern kernel: ${\bm{k}}\in\{0,1\}^{r\times r}$; Stride: $s=1$; Max iteration number: $T$; Tolerance: $t$; group $l_{0}$ budget: $\epsilon$; Step size: $\alpha$, $\beta$; Small constant: $\gamma=2\times 10^{-8}$
2: Random initialize $\bm{p}$} and $\widetilde{{\bm{v}}}$
[…]
5:for i=0,1,,T1i=0,1,...,T-1 do
6:  Calculate the loss $\mathcal{L}({\bm{\theta}},{\bm{x}}+{\bm{p}}\odot{\bm{m}})$
7:  if unprojected then
8:   $g_{{\bm{p}}}=\nabla_{{\bm{\delta}}}{\mathcal{L}}\odot\min\left(\mathrm{TConv}(\sigma(\widetilde{{\bm{v}}}),{\bm{k}},s),1\right)$
9:  else
10:   $g_{{\bm{p}}}=\nabla_{\bm{\delta}}{\mathcal{L}}\odot{\bm{m}}$
11:  end if
[…]
13:  ${\bm{p}}=\Pi_{{\mathcal{S}}_{{\bm{p}}}}({\bm{p}}+\alpha\cdot\mathtt{sign}(g_{{\bm{p}}}))$
14:  ${\bm{d}}=g_{\widetilde{{\bm{v}}}}/\|g_{\widetilde{{\bm{v}}}}\|_{2}$ if $g_{\widetilde{{\bm{v}}}}||_{2}\geq\gamma$ else 0
15:  ${\bm{v}}_{old},\ \widetilde{{\bm{v}}}={\bm{v}},\ \widetilde{{\bm{v}}}+\beta\cdot{\bm{d}}$
16:  $){\bm{v}}=\Pi_{{\mathcal{S}}_{{\bm{v}}}}(\sigma(\widetilde{{\bm{v}}}$))
17:  ${\bm{m}}=\min\left(\mathrm{TConv}({\bm{v}},{\bm{k}},s),1\right)$
18:  if attack succeeds then
19:   break
20:  end if
21:  if ${\bm{v}}-{\bm{v}}_{old}||_{0}\leq 0$ for $t$ consecutive iters then
22:   Random initialize $\widetilde{{\bm{v}}}$
23:  end if
24:end for
25:Output: Perturbation: ${\bm{\delta}}={\bm{p}}\odot{\bm{m}}$
IV-C sAA for Structured Sparse Perturbations
Similar to Section III-B, we also combine sPGDunproj, sPGDproj and Sparse-RS to build Sparse-AutoAttack (sAA) for reliable evaluation of robustness against structured sparse perturbations. We extended the original version of Sparse-RS so that it can generate structured sparse perturbations with any customized patterns. Specifically, we just apply a pattern mask to the patches generated by Sparse-RS.
V Adversarial Training
In addition to adversarial attacks, we explore adversarial training to build robust models against sparse perturbations. In the framework of adversarial training, the attack is used to generate adversarial perturbation in each training iteration, so the attack algorithm should not be too computationally expensive. In this regard, we run the untargeted sPGD (Algorithm 1) for $20$ iterations to generate sparse adversarial perturbations during training. We incorporate sPGD in the framework of vanilla adversarial training [4] and TRADES [25] and name corresponding methods sAT and sTRADES, respectively. Note that we use sAT and sTRADES as two examples of applying sPGD to adversarial training since sPGD can be incorporated into any other adversarial training variant as well. To accommodate the scenario of adversarial training, we make the following modifications to sPGD.
Random Backward Function: Since the sparse gradient and the unprojected gradient as described in Section III-A induce different exploration-exploitation trade-offs, we randomly select one of them to generate adversarial perturbations for each mini-batch when using sPGD to generate adversarial perturbations. Compared with mixing these two backward functions together, as in sAA, random backward function does not introduce computational overhead.
Multi-$\epsilon$ Strategy: Inspired by $l_{1}$-APGD [19] and Fast-EG-$l_{1}$ [20], multi-$\epsilon$ strategy is adopted to strengthen the robustness of model. That is, we use a larger sparsity threshold, i.e., $\epsilon$ in Algorithm 1, in the training phase than in the test phase.
Higher Tolerance for Reinitialization: The default tolerance for reinitialization in sPGD is $3$ iterations, which introduces strong stochasticity. However, in the realm of adversarial training, we have a limited number of iterations. As a result, the attacker should focus more on the exploitation ability to ensure the strength of the generated adversarial perturbations. While stochasticity introduced by frequent reinitialization hurts exploitation, we find a higher tolerance for reinitialization improves the performance. In practice, we set the tolerance to $10$ iterations in adversarial training.
VI Experiments
In this section, we conduct extensive experiments to compare our attack methods with baselines in evaluating the robustness of various models against $l_{0}$ bounded and structured sparse perturbations. Besides the effectiveness with an abundant query budget, we also study the efficiency of our methods. Our results demonstrate that our sPGD performs best among white-box attacks. With limited iterations, sPGD achieves significantly better performance than existing methods. Therefore, sAA, consisting of the best white-box and black-box attacks, has the best attack success rate. sPGD, due to its efficiency, is utilized for adversarial training to obtain the best robust models against $l_{0}$ bounded and structured sparse adversarial perturbations. To further demonstrate the efficacy of our method, we evaluate the transferability of sPGD. The results show that sPGD has a high transfer success rate, making it applicable in practical scenarios. Moreover, to exhibit the versatility of our method, we evaluate sPGD in objection detection and segmentation tasks. The adversarial examples generated by our methods are presented in Sec. VI-H. In addition, we conduct ablation studies for analysis in Appendix C. Implementation details are deferred to Appendix A.
TABLE I: Robust accuracy of various models on different sparse attacks. The models are trained on CIFAR-10. Note that we report results of Sparse-RS (RS) with fine-tuned hyperparameters, which outperforms its original version in Croce et al. [17]. CornerSearch (CS) is evaluated on 1000 samples due to its high computational complexity.
Model Network Clean Black White sAA
CS RS SF PGD0 SAIF sPGDp sPGDu
Vanilla RN-18 93.9 1.6 0.0 25.3 2.1 12.0 0.0 0.0 0.0
$l_{\infty}$-adv. trained, $\epsilon=8/255$
GD PRN-18 87.4 30.5 12.2 61.1 36.0 51.3 17.1 24.3 11.3
PORT RN-18 84.6 30.8 15.2 62.1 31.4 52.1 17.4 23.0 13.0
DKL WRN-28 92.2 35.3 13.2 62.5 41.2 52.3 18.4 24.9 12.1
DM WRN-28 92.4 34.8 12.6 57.9 38.5 49.4 17.9 24.0 11.6
$l_{2}$-adv. trained, $\epsilon=0.5$
HAT PRN-18 90.6 38.9 23.5 65.3 35.4 60.2 19.0 18.6 16.6
PORT RN-18 89.8 36.8 20.6 64.3 30.6 59.7 16.0 15.5 13.8
DM WRN-28 95.2 48.5 27.7 68.2 47.5 70.9 25.0 26.7 22.1
FDA WRN-28 91.8 47.8 31.1 71.8 40.1 68.2 28.0 31.4 25.5
$l_{1}$-adv. trained, $\epsilon=12$
$l_{1}$-APGD PRN-18 80.7 41.3 36.5 70.3 50.5 62.3 30.4 31.3 29.0
Fast-EG-$l_{1}$ PRN-18 76.2 40.7 34.8 64.9 46.7 56.9 29.6 30.1 28.0
$l_{0}$-adv. trained, $\epsilon=15$
PGD0-A PRN-18 83.7 17.5 6.1 73.7 62.9 60.5 19.4 27.5 5.6
PGD0-T PRN-18 90.5 19.5 7.2 85.5 63.6 69.8 31.4 41.2 7.1
sAT PRN-18 80.9 46.0 37.6 77.1 74.1 72.3 71.2 70.3 37.6
sTRADES PRN-18 90.3 71.7 63.7 89.5 88.1 86.5 85.9 83.8 63.7
(a) $\epsilon=$ 15
Model Network Clean Black White sAA
CS RS SF PGD0 SAIF sPGDp sPGDu
Vanilla RN-18 93.9 1.2 0.0 17.5 0.4 3.2 0.0 0.0 0.0
$l_{\infty}$-adv. trained, $\epsilon=$ 8/255
GD PRN-18 87.4 26.7 6.1 52.6 25.2 40.4 9.0 15.6 5.3
PORT RN-18 84.6 27.8 8.5 54.5 21.4 42.7 9.1 14.6 6.7
DKL WRN-28 92.2 33.1 7.0 54.0 29.3 41.1 9.9 15.8 6.1
DM WRN-28 92.4 32.6 6.7 49.4 26.9 38.5 9.9 15.1 5.9
$l_{2}$-adv. trained, $\epsilon=$ 0.5
HAT PRN-18 90.6 34.5 12.7 56.3 22.5 49.5 9.1 8.5 7.2
PORT RN-18 89.8 30.4 10.5 55.0 17.2 48.0 6.3 5.8 4.9
DM WRN-28 95.2 43.3 14.9 59.2 31.8 59.6 13.5 12.0 10.2
FDA WRN-28 91.8 43.8 18.8 64.2 25.5 57.3 15.8 19.2 14.1
$l_{1}$-adv. trained, $\epsilon=$ 12
$l_{1}$-APGD PRN-18 80.7 32.3 25.0 65.4 39.8 55.6 17.9 18.8 16.9
Fast-EG-$l_{1}$ PRN-18 76.2 35.0 24.6 60.8 37.1 50.0 18.1 18.6 16.8
$l_{0}$-adv. trained, $\epsilon=$ 20
PGD0-A PRN-18 77.5 16.5 2.9 62.8 56.0 47.9 9.9 21.6 2.4
PGD0-T PRN-18 90.0 24.1 4.9 85.1 61.1 67.9 27.3 37.9 4.5
sAT PRN-18 84.5 52.1 36.2 81.2 78.0 76.6 75.9 75.3 36.2
sTRADES PRN-18 89.8 69.9 61.8 88.3 86.1 84.9 84.6 81.7 61.7
(b) $\epsilon=$ 20
VI-A Robustness against Unstructured Sparse Perturbations
First, we compare our proposed sPGD, including sPGDproj (sPGDp) and sPGDunproj (sPGDu) as defined in Section III-B, and sAA with existing white-box and black-box attacks that generate $l_{0}$ bounded sparse perturbations. We evaluate different attack methods based on the models trained on CIFAR-10 [64] and report the robust accuracy with $\epsilon=15,~20$ on the whole test set in Table II(b). Additionally, the results on ImageNet-100 [65] and a real-world traffic sign dataset GTSRB [66] are reported in Table II. Note that the image sizes in GTSRB vary from $15\times 15$ to $250\times 250$. For convenience, we resize them to $224\times 224$ and use the same model architecture as in ImageNet-100. Furthermore, only the training set of GTSRB has annotations, we manually split the original training set into a test set containing $1000$ instances and a new training set containing the rest data. In Appendix B, we report more results on CIFAR-10 and CIFAR-100 [64] in Table VIII, to demonstrate the efficacy of our methods.
Models: We select various models to comprehensively evaluate their robustness against $l_{0}$ bounded perturbations. As a baseline, we train a ResNet-18 (RN-18) [67] model on clean inputs. For adversarially trained models, we select competitive models that are publicly available, including those trained against $l_{\infty}$, $l_{2}$ and $l_{1}$ bounded perturbations. For the $l_{\infty}$ case, we include adversarial training with the generated data (GD) [46], the proxy distributions (PORT) [44], the decoupled KL divergence loss (DKL) [48] and strong diffusion models (DM) [49]. For the $l_{2}$ case, we include adversarial training with the proxy distributions (PORT) [44], strong diffusion models (DM) [49], helper examples (HAT) [47] and strong data augmentations (FDA) [45]. The $l_{1}$ case is less explored in the literature, so we only include $l_{1}$-APGD adversarial training [19] and the efficient Fast-EG-$l_{1}$ [20] for comparison. The network architecture used in these baselines is either ResNet-18 (RN-18), PreActResNet-18 (PRN-18) [67] or WideResNet-28-10 (WRN-28) [68]. For the $l_{0}$ case, we evaluate PGD0 [13] in vanilla adversarial training (PGD0-A) and TRADES (PGD0-T) using the same hyper-parameter settings as in Croce and Hein [13]. Since other white-box sparse attacks present trivial performance in adversarial training, we do not include their results. Finally, we use our proposed sPGD in vanilla adversarial training (sAT) and TRADES (sTRADES) to obtain PRN-18 models to compare with these baselines.
Attacks: We compare our methods with various existing black-box and white-box attacks that generate $l_{0}$ bounded perturbations. The black-box attacks include CornerSearch (CS) [13] and Sparse-RS (RS) [17]. The white-box attacks include SparseFool (SF) [12], PGD0 [13] and Sparse Adversarial and Interpretable Attack Framework (SAIF) [41]. The implementation details of each attack are deferred to Appendix A. Specifically, to exploit the strength of these attacks in reasonable running time, we run all these attacks for either $10000$ iterations or the number of iterations where their performances converge. Note that, the number of iterations for all these attacks are no smaller than their default settings. In addition, we report the results of RS with fine-tuned hyperparameters, which outperforms its default settings in [17]. Finally, we report the robust accuracy under CS attack based on only $1000$ random test instances due to its prohibitively high computational complexity.
Based on the results in Table II(b) and Table II, we can find that SF attack, PGD0 attack and SAIF attack perform significantly worse than our methods for all the models studied. That is, our proposed sPGD always performs the best among white-box attacks. Among black-box attacks, CS attack can achieve competitive performance, but it runs dozens of times longer than our method does. Therefore, we focus on comparing our method with RS attack. For $l_{1}$ and $l_{2}$ models, our proposed sPGD significantly outperforms RS attack. By contrast, RS attack outperforms sPGD for $l_{\infty}$ and $l_{0}$ models. This gradient masking phenomenon is, in fact, prevalent across sparse attacks. Given sufficient iterations, RS outperforms all other existing white-box attacks for $l_{\infty}$ and $l_{0}$ models. Nevertheless, among white-box attacks, sPGD exhibits the least susceptibility to gradient masking and has the best performance. The occurrence of gradient masking in the context of $l_{0}$ bounded perturbations can be attributed to the non-convex nature of adversarial budgets. In practice, the perturbation updates often significantly deviate from the direction of the gradients because of the projection to the non-convex set. Similar to AA in the $l_{1}$, $l_{2}$ and $l_{\infty}$ cases, sAA consists of both white-box and black-box attacks for comprehensive robustness evaluation. It achieves the best performance in all cases in Table II(b) and Table II by a considerable margin.
TABLE II: Robust accuracy of various models on different sparse attacks. Our sAT model is trained with $\epsilon=1200$. (a) Results on ImageNet-100 [65]. (b) Results on GTSRB [66]. Note that the results of Sparse-RS (RS) with tuned hyperparameters are reported. All models are RN-34, and are evaluated on $500$ samples. The results of SparseFool (SF) and PGD0 are included due to their poor performance, and CornerSearch (CS) is not evaluated here due to its high computational complexity, i.e. nearly 1 week on one GPU for each run.
Model Clean Black White sAA
RS SAIF sPGDp sPGDu
Vanilla 83.0 0.2 0.6 0.2 0.4 0.0
$l_{0}$-adv. trained, $\epsilon=200$
PGD0-A 76.0 6.8 11.0 1.8 18.8 1.8
sAT 86.2 61.4 69.0 78.0 77.8 61.2
(a) ImageNet-100, $\epsilon=$ 200
Model Clean Black White sAA
RS SAIF sPGDp sPGDu
Vanilla 99.9 18.0 9.5 0.3 0.3 0.3
$l_{0}$-adv. trained, $\epsilon=600$
PGD0-A 99.8 37.6 13.8 0.0 0.0 0.0
sAT 99.8 88.4 96.2 88.6 96.2 85.4
(b) GTSRB, $\epsilon=$ 600
In the case of $l_{0}$ adversarial training, the models are adversarially trained against sparse attacks. However, Figure 3 illustrates that the performance of RS attack on standard, $l_{1}$-, $l_{2}$- and $l_{\infty}$-trained models drastically deteriorates with limited iterations (e.g., smaller than $100$), so RS is not suitable for adversarial training where we need to generate strong adversarial perturbations in limited number iterations. Empirical evidence suggests that employing RS with $20$ iterations for adversarial training, i.e., the same number of iterations as in other methods, yields trivial performance, so it is not included in Table II(b) or Table II for comparison. In addition, models trained by PGD0-A and PGD0-T, which generate $l_{0}$ bounded perturbations, exhibit poor robustness to various attack methods, especially sAA. By contrast, the models trained by sAT and sTRADES show the strongest robustness, indicated by the comprehensive sAA method and all other attack methods. Compared with sAT, sTRADES achieves better performance in both robustness and accuracy. Finally, models trained with $l_{1}$ bounded perturbations are the most robust ones among existing non-$l_{0}$ training methods. It could be attributed to the fact that $l_{1}$ norm is the tightest convex relaxation of $l_{0}$ norm [69]. From a qualitative perspective, $l_{1}$ attacks also generate relatively sparse perturbations [20], which makes the corresponding model robust to sparse perturbations to some degree.
Our results indicate sPGD and RS can complement each other. Therefore, sAA, an AutoAttack-style attack that ensembles both attacks achieves the state-of-the-art performance on all models. It is designed to have a similar computational complexity to AutoAttack in $l_{\infty}$, $l_{2}$ and $l_{1}$ cases.
VI-B Robustness against Structured Sparse Perturbations
TABLE III: Robust accuracy of various models on different structured sparse attacks. (a) Results on CIFAR-10, all models are PRN-18. (b) Results on ImageNet-100, the test set contains $500$ samples, all models are RN-34. Note that the evaluated group sparsity levels are approximately equivalent to the $l_{0}$ sparsity levels used in Table II(b) and II, and the Sparse-RS (RS) used here is the proposed extended version.
Model Clean Black White sAA
RS LOAP sPGDp sPGDu
row, $\epsilon=1$
Vanilla 93.9 22.5 - 1.0 1.3 1.0
sTRADES-$l_{0}$ 89.8 83.1 - 54.6 65.0 53.9
sTRADES-row 89.3 85.9 - 77.3 83.7 77.3
$3\times 3$ patch, $\epsilon=2$
Vanilla 93.9 6.7 4.5 1.9 6.6 1.7
LOAP+TRADES 92.8 75.6 66.4 47.4 71.7 47.1
sTRADES-$l_{0}$ 89.8 65.5 60.0 46.9 67.5 46.7
sTRADES-p3x3 87.9 77.5 81.4 72.8 81.7 72.3
$5\times 5$ patch, $\epsilon=1$
Vanilla 93.9 7.5 2.3 1.7 3.2 1.7
LOAP+TRADES 92.4 76.9 61.7 37.0 60.8 36.0
sTRADES-$l_{0}$ 89.8 56.4 35.0 26.7 44.3 26.7
sTRADES-p5x5 89.5 72.9 73.1 56.0 73.3 55.9
(a) CIFAR-10
Model Clean Black White-Box sAA
RS LOAP sPGDp sPGDu
row, $\epsilon=1$
Vanilla 83.0 54.8 - 5.2 8.0 5.2
sAT-$l_{0}$ 86.2 78.6 - 65.8 73.8 65.8
sAT-row 81.0 78.2 - 76.6 78.4 76.6
$10\times 10$ patch, $\epsilon=2$
Vanilla 83.0 13.0 4.6 4.6 6.4 4.0
LOAP+AT 83.8 57.0 72.4 0.0 0.4 0.0
sAT-$l_{0}$ 86.2 27.8 10.8 10.6 33.0 10.0
sAT-p10x10 81.6 64.0 62.8 35.8 73.2 35.8
$14\times 14$ patch, $\epsilon=1$
Vanilla 83.0 7.5 2.3 1.7 12.0 1.7
LOAP+AT 83.4 53.0 20.8 0.0 530 0.0
sAT-$l_{0}$ 86.2 28.3 10.2 15.4 34.8 14.6
sAT-p14x14 80.0 57.2 73.0 41.2 71.2 39.6
(b) ImageNet-100
Apart from unstructured sparse perturbations, we evaluate the effectiveness of our method in Algorithm 2 in generating structured sparse perturbations in this subsection. Given the paucity of available methods for comparison, we focus on comparing our method with the extended Sparse-RS (RS) and LOAP [16], which is a white-box approach generating adversarial patches, on CIFAR-10 and ImageNet-100. For a comprehensive evaluation, we include the results of different types of structured sparse perturbations, e.g., row and patches with different group sparsity level and different sizes. It should be noted that the unstructured $l_{0}$ norms of the structured perturbations studied here are approximately the same as the perturbations in Table II(b) and II. In addition, we evaluate the attacks on vanilla models, models trained with unstructured $l_{0}$ bounded perturbations, and those trained with the specific structured sparse perturbations.
The results in Table III indicate that the proposed sAA also achieves the state-of-the-art performance in generating structured sparse perturbations in almost all cases. Furthermore, the models trained with our methods exhibit the strongest robustness against structured sparse perturbations. As anticipated, models trained on specific structured sparse adversarial examples significantly outperform those trained on unstructured sparse adversarial examples. This highlights the inherent limitation of adversarially trained models in maintaining robustness against unseen types of perturbations during training.
VI-C Adversarial Watermarks
Although only a few pixels are perturbed in sparse attacks, such perturbations are still perceptible due to their unconstrained magnitude. In this subsection, we additionally constrain the $l_{\infty}$ norm of magnitude $\bm{p}$}, meaning that the feasible set for $vp$ is rewritten as ${\mathcal{S}}_{\bm{p}}=\{{\bm{p}}\in\mathbb{R}^{h\times w\times c}|\|{\bm{p}}\|_{\infty}\leq\epsilon_{\infty},~0\leq{\bm{x}}+{\bm{p}}\leq 1\}$. Combining with structured sparsity constraint, we can generate alleged adversarial watermarks with our method.
TABLE IV: Robust accuracy of vanilla model (RN-34) against adversarial watermarks with different $l_{\infty}$ adversarial budgets. The experiment is conducted on ImageNet-100.
$\epsilon_{\infty}$ $8/255$ $16/255$ $32/255$ $64/255$
RS 59.0 44.6 32.4 26.4
sPGDp 47.2 25.4 16.6 9.0
sPGDu 56.4 39.6 25.8 16.8
sAA 44.2 24.0 15.6 8.8
(a) $\times18$ circle
$\epsilon_{\infty}$ $8/255$ $16/255$ $32/255$ $64/255$
RS 14.2 12.8 9.2 5.8
sPGDp 3.8 2.6 2.4 2.2
sPGDu 7.4 4.8 2.8 2.0
sAA 2.0 0.8 0.8 0.6
(b) $\times60$ letter “A”
[…]
VI-D Comparison under Different Iteration Numbers and Different Sparsity Levels
(a) CIFAR-10, $\epsilon=20$
(b) ImageNet-100, $\epsilon=200$
(c) CIFAR-10, $t=10000$
Figure 3: Comparison between sPGD and RS attack under different iteration numbers and different sparsity levels. (a) Different iteration number comparison on CIFAR-10, $\epsilon=20$. ResNet18 (std), PORT ($l_{\infty}$ and $l_{2}$) [44], $l_{1}$-APGD ($l_{1}$) [19] and sTRADES ($l_{0}$) are evaluated. (b) Different iteration number comparison on ImageNet-100, $\epsilon=200$. ResNet34 (std), Fast-EG-$l_{1}$ ($l_{1}$) [20] and sAT ($l_{0}$) are evaluated. In (a) and (b), the total iteration number ranges from 2020 and $10000$. For better visualization, the x-axis is in the log scale. (c) Different sparsity comparison on CIFAR-10. The evaluated models are the same as those in (a). The $\epsilon$ ranges from 0 and $50$. The number of total iterations is set to $10000$. Note that the results of sPGD and RS attack are shown in solid lines and dotted lines, respectively.
In this subsection, we further compare our method sPGD, which is a white-box attack, with RS attack, the strongest black-box attack in the previous section. Specifically, we compare them under various iteration numbers on CIFAR-10 and ImageNet-100, which have different resolutions. In addition, we also compare sPGD and RS under different sparsity levels on CIFAR-10.
As illustrated in Figure 3 (a) and (b), sTRADES has better performance than other robust models by a large margin in all iterations of both sPGD and RS attacks, which is consistent with the results in Table II(b)$I$ and VIII. For vanilla and other robust models, although the performances of both sPGD and RS attack get improved with more iterations, sPGD significantly outperforms RS attack when the iteration number is small (e.g. <1000<1000 iterations), which makes it feasible for adversarial training. Similar to other black-box attacks, the performance of RS attack drastically deteriorates when the query budget is limited. In addition, our proposed gradient-based sPGD significantly outperforms RS on ImageNet-100, where the search space is much larger than that on CIFAR-10, i.e., higher image resolution and higher sparsity level $\epsilon$. This suggests that our approach is scalable and shows higher efficiency on high-resolution images. Furthermore, although the performance of RS does not converge even when the iteration number reaches $10000$, a larger query budget will make it computationally impractical. Following the setting in Croce et al. [17], we do not consider a larger query budget in our experiments, either.
Furthermore, we can observe from Figure 3 (c) that RS attack shows slightly better performance only on the $l_{\infty}$ model and when $\epsilon$ is small. The search space for the perturbed features is relatively small when $\epsilon$ is small, which facilitates heuristic black-box search methods like RS attack. As $\epsilon$ increases, sPGD outperforms RS attack in all cases until both attacks achieve almost $100\%$ attack success rate.
VI-E Efficiency of sPGD
To further showcase the efficiency of our approach, we first compare the distributions of the number of iterations needed by sPGD and RS to successfully generate adversarial samples. Figure 4 illustrates the distribution of the iteration numbers needed by sPGD and RS to successfully generate adversarial samples. For $l_{\infty}$ and $l_{1}$ robust models, our proposed sPGD consumes distinctly fewer iteration numbers to successfully generate an adversarial sample than RS, the strongest black-box attack in Table II(b), while maintaining a high attack success rate. Similar to the observations in Figure 3, the model trained by sTRADES suffers from gradient masking. However, RS still requires a large query budget to successfully generate an adversarial sample. This further demonstrates the efficiency of sPGD, which makes adversarial training feasible.
(a) $l_{\infty}$ PORT [44]
(b) $l_{1}$-APGD [19]
(c) sTRADES
Figure 4: Distribution of the iteration numbers needed by sPGD (blue) and RS (orange) to successfully generate adversarial samples. The results are obtained from different models: (a) $l_{\infty}$ PORT [44], (b) $l_{1}$-APGD [19] and (c) our sPGD. The average iteration numbers (Avg.) and attack success rate (ASR), i.e., 11-Robust Acc., are reported in the legend. For better visualization, we clip the minimum iteration number to $10$ and show the x- and y-axis in log scale.
Additionally, we compare the runtime of sPGD with other attacks in the same configuration as in Table II(b). As shown in Table V, the proposed sPGD shows the highest efficiency among various attacks. Although sAA consumes more time (approximately $2\times$ sPGD + RS), it can provide a reliable evaluation against $l_{0}$ bounded perturbation.
TABLE V: Runtime of different attacks on 1000 test instances with batch size 500. The sparsity level $\epsilon=20$. The evaluated model is sTRADES. The model is trained on CIFAR-10. The experiments are implemented on NVIDIA Tesla V100.
Attack CS RS SF PGD0
Runtime 604 min 59 min 92 min 750 min
Attack SAIF sPGDp sPGDu sAA
Runtime 122 min 42 min 45 min 148 min
VI-F Transferability of Adversarial Perturbations
TABLE VI: Transferability of RS and sPGD between VGG11 (V) and ResNet18 (R) on CIFAR-10 with $\epsilon=$ 20. Attack success rate (ASR) is reported. The perturbations are generated on the source model (left), and are evaluated on the target model (right). Note that vanilla models are evaluated, and $\alpha$ and $\beta$ of sPGD are set to 0.75.
$\rightarrowV V\rightarrowR$ $\rightarrowR R\rightarrowV$
RS 53.9 28.4 33.0 37.4
sPGDp 58.0 43.9 50.8 48.0
sPGDu 64.9 40.0 52.7 56.7
To evaluate the transferability of our attack across different models and architectures, we generate adversarial perturbations based on one model and report the attack success rate (ASR) on another model. As shown in Table VI, sPGD exhibits better transferability than the most competitive baseline, Sparse-RS (RS) [17]. This further demonstrates the effectiveness of our method and its potential application in more practical scenarios.
VI-G Attack Object Detection and Segmentation Models
The versatility of our method allows it to be applied to a variety of tasks beyond classification, e.g., object detection and segmentation. For implementation, one could simply replace the loss function $\mathcal{L}$ in Algorithms 1 and 2 with a specific detection or segmentation loss function.
In Table VII, we report the mAP50-95 of YOLOv8-M [70] in object detection and segmentation tasks when the input image is manipulated by sparse perturbations, including both structured and unstructured ones. In addition, we include the results of RS and SAIF as a comparison. Since SAIF is not able to generate structured sparse perturbations, we do not include its results in the experiments of generating patch perturbations. The results indicate that sPGD, as an attack framework, is still effective in these two tasks. Considering higher resolution images in the COCO dataset [71] ($640\times 640$), we use approximately the same ratio between the number of perturbed pixels and the total number of pixels as in Table II when setting the perturbation budgets $\epsilon$, i.e., $0.4\%$ / $0.8\%$ of the pixels. Compared with structured perturbations, unstructured perturbations are spatially more dispersed, making them more effective for multi-object images in these tasks.
TABLE VII: Evaluation of YOLOv8-M performance on clean and adversarial COCO dataset [71] in object detection and segmentation tasks. The metric is mAP50-95. We test the performance on the adversarial examples generated by different attacks, where both unstructured and structured cases are considered.
Mode $\epsilon$ Clean RS SAIF sPGD
$l_{0}$ 1600 55.1 17.1 15.8 15.5
3200 55.1 8.1 7.4 6.6
$4\times 4$ patch 100 55.1 47.1 - 40.1
200 55.1 40.8 - 34.5
(a) Object Detection
Mode $\epsilon$ Clean RS SAIF sPGD
$l_{0}$ 1600 53.6 17.6 16.4 15.6
3200 53.6 8.8 8.2 7.4
$4\times 4$ patch 100 53.6 46.5 - 39.1
200 53.6 39.7 - 33.1
(b) Segmentation
VI-H Visualization
(a) $l_{0}$, $\epsilon=200$
(b) $l_{0}$, $\epsilon=200$
(c) row, $\epsilon=1$
(d) row, $\epsilon=1$
(e) star, $25\times 25$, $\epsilon=1$
(f) star, $25\times 25$, $\epsilon=1$
(g) “A”, $60\times 60$, $\epsilon_{\infty}=32/255$
(h) “A”, $60\times 60$, $\epsilon_{\infty}=32/255$
[…]
For a better visualization, we present some adversarial examples with different types of sparse perturbations in Figure 5, including $l_{0}$ bounded perturbations, row-wise perturbations, star-like perturbations, and adversarial watermarks of letter “A”. The attack is sPGD, and the model is vanilla ResNet-34 trained on ImageNet-100. More adversarial examples are shown in Appendix D.
VII Conclusion
In this paper, we propose an effective and efficient white-box attack named sPGD to generate both unstructured and structured perturbations. sPGD obtains the state-of-the-art performance among white-box attacks. Based on this, we combine it with black-box attacks for more comprehensive and reliable evaluation of robustness against sparse perturbations. Our proposed sPGD is particularly effective in the realm of limited iteration numbers. Due to its efficiency, we incorporate sPGD into the framework of adversarial training to obtain robust models against sparse perturbations. The models trained with our method demonstrate the best robust accuracy.
Acknowledgments
This work is supported by the NSFC project (No. 62306250). It is supported by the internal funds of City University of Hong Kong (No. 9610614 and No. 9229130). We also thank Yixiao Huang for his contribution to this work.